OpenStudy (juscallmesteve):

Using l'Hopitals rule to find Lim x-->0 sin5x/sinx, I will post my attempt inside would like someone to check

OpenStudy (juscallmesteve):

\[\lim_{x \rightarrow 0} \frac{ \sin5x }{ sinx } == 5\]

OpenStudy (juscallmesteve):

so first i took \[\lim_{x \rightarrow 0} \frac{ \sin5x }{ sinx }\] and then took the \[\frac{ d }{ dx }\] of top and bottom and got \[\lim_{x \rightarrow 0} \frac{ 5\cos5x }{ cosx }\] then evaluated the equation using the limit

OpenStudy (welshfella):

looks good

OpenStudy (juscallmesteve):

it came to 5 *(cos(infinity)) which is 5*1 which is 5 and cos(infinity) = 1 so 5/1 = 5

OpenStudy (juscallmesteve):

Thank you

OpenStudy (welshfella):

yep you only needed one application of lhopitals tis time.

OpenStudy (juscallmesteve):

awesome thank you again

OpenStudy (welshfella):

yw